# Observational Constraints on Dark Energy Models with $\Lambda$ as an Equilibrium Point

### v1

@@ Abstract
We investigate a dynamical reconstruction of the dark energy equation of state parameter by assuming that it satisfies a law of motion described by an autonomous second-order differential equation, with the limit of the cosmological constant as an equilibrium point. We determine the asymptotic solutions of this equation and use them to construct two families of parametric dark energy models, employing both linear and logarithmic parametrizations with respect to the scale factor. We perform observational constraints by using the Supernova, the Cosmic Chronometers and the Baryon Acoustic Oscillations. The constraint parameters are directly related with the initial value problem for the law of motion and its algebraic properties. The analysis shows that most of the models fit the observational data well with a preference to the models of the logarithmic parametrization. Furthermore, we introduce a new class of models as generalizations of the CPL model, for which the equilibrium point is a constant value rather than the cosmological constant. These models fit the data in a similar or better way to the CPL and the $\Lambda$CDM cosmological models.
@@ I Introduction
The determination of the equation of state parameter for the dark energy component can provide crucial insights into the nature of this exotic fluid source. Dynamical dark energy models with a time-varying equation of state parameter $w\left(a\right)$ have been introduced to explain cosmological observations. The Chevallier, Polarski and Linder cpl1 ; cpl1a (CPL) parametrization is the simplest varying dark energy model, introducing two degrees of freedom. The CPL model helps to determine how distinct a given dark energy model is from the cosmological constant cpl2 ; cpl3 ; cpl4 ; clp5 . Due to the limitations of the CPL model, various alternative dark energy models have been proposed over the past two decades. For more details, we refer the reader to par1 ; par2 ; par4 ; par5 ; par6 ; par7 ; exp1 ; exp2 ; osc1 ; osc2 ; osc3 ; osc4 ; osc5 and the references therein.
@@ II Cosmological Framework
$\left(\dot{\rho}_{r}+4H\rho_{r}\right)+\dot{\rho}_{b}+\dot{\rho}_{m}+\left(\dot{\rho}_{d}+3\left(1+w_{d}\right)H\rho_{d}\right)=0.$ (13)
@@
$\rho_{r}=\rho_{r,0}\left(\frac{a}{a_{0}}\right)^{-4},~{}\rho_{b}=\rho_{b,0}\left(\frac{a}{a_{0}}\right)^{-3},~{}\rho_{m}=\rho_{m,0}\left(\frac{a}{a_{0}}\right)^{-3}$ (14)
@@
$\Omega_{r0}+\Omega_{b0}+\Omega_{m0}+\Omega_{d0}=1.$ (17)
@@
$H\left(a\right)=H_{0}\sqrt{\Omega_{r0}\left(\frac{a}{a_{0}}\right)^{-3}+\Omega_{b0}\left(\frac{a}{a_{0}}\right)^{-3}+\Omega_{m0}\left(\frac{a}{a_{0}}\right)^{-3}+\Omega_{\Lambda 0}}.$ (18)
@@
$H\left(a\right)=H_{0}\sqrt{\Omega_{r0}\left(\frac{a}{a_{0}}\right)^{-3}+\Omega_{b0}\left(\frac{a}{a_{0}}\right)^{-3}+\Omega_{m0}\left(\frac{a}{a_{0}}\right)^{-3}+\Omega_{d0}a^{-3\left(1+w_{0}\right)}}.$ (19)
@@ III Dynamical dark energy with $\Lambda$ as an Equilibrium Point
where now $w_{a}=\operatorname{Re}\left(\Delta_{+}\right)=\operatorname{Re}\left(\Delta_{-}\right)$ and $w_{b}=\operatorname{Im}\left(\Delta_{+}\right)=\operatorname{Im}\left(\Delta_{-}\right)$.
@@ III.3 Parametric space
$w_{+}^{I}\left(z\right)$ $w_{a}$ $-1+e^{w_{a}Z}$
$w_{-}^{I}\left(z\right)$ $w_{a}$ $-1-e^{w_{a}Z}$
@@ IV Observational Data
The Pantheon+ dataset [^1] https://github.com/PantheonPlusSH0ES/DataRelease comprises 1701 light curves of 1550 spectroscopically confirmed supernova events within the range $2.2710^{-3}<z<$ 2.27. The data provides the distance modulus $\mu^{obs}~{}$at observed redshifts $z$. The theoretical distance modulus is constructed as follows $\mu^{th}=5\log D_{L}+25,$ where $D_{L}$ is the luminosity distance defined by the Hubble functions from the expression $D_{L}=c\left(1+z\right)\int\frac{dz}{H\left(z\right)}$.
For the Cosmic Chronometers we use 31 direct measurements of the Hubble parameter across redshifts in the range [^2] https://github.com/Ahmadmehrabi/Cosmic_chronometer_data $0.09\leq z\leq 1.965~{}$ as summarized in Table 1 in sunny .
The BAO data are provided by the SDSS Galaxy Consensus, quasars and Lyman-$\alpha$ forests bbn0 and the data from the DESI collaboration des1 ; des2 ; des3 . These data provide observation values of the $\frac{D_{M}}{r_{s}},~{}\frac{D_{V}}{r_{s}}$, $\frac{D_{H}}{r_{s}}$ where $r_{s}$ is the sound horizon at the drag epoch, $D_{M}$ is the comoving angular distance $D_{M}=\left(1+z\right)^{-1}D_{L}$ ; $D_{V}$ is the volume averaged distance $D_{V}^{3}=cD_{L}\frac{z}{H\left(z\right)}$ and $D_{H}$ is the Hubble distance $D_{H}=\frac{c}{H}$.
@@ IV.1 Methodology
In order to obtain the optimal values for the cosmological parameters we employ the Bayesian inference COBAYA [^3] https://cobaya.readthedocs.io/ cob1 ; cob2 with a custom theory and the MCMC sampler. From the results of the Planck 2018 collaborations we consider the energy density for the radiation to be $\Omega_{r0}=4.15~{}10^{-5}$, and for the baryons $\Omega_{b0}=0.0486$. Thus, we constrain our proposed models for the parameters $\left\{H_{0},\Omega_{m0},\mathbf{w,}r_{drag}\right\},~{}$where $\mathbf{w}$ describes the vector of parameters within the equation of state parameters, as they are given in Table 1.
For all the runs we selected the that $H_{0}\in\left[60,75\right]$,${}\Omega_{m0}\in\left[0.15,0.35\right]$ and $r_{drag}\in\left[130,160\right]$.
@@ V Results
We constrain the dynamical dark energy models using two sets of data: (I) Supernova and Cosmic Chronometers ($SN+CC$) and (II) Supernova, Cosmic Chronometers and Baryonic Acoustic Oscillations ($SN+CC+BAO$). We determine the free parameters along with their corresponding $1\sigma$ errors by locating the maximum of the likelihood functions. The results for the two tests are presented in Tables 2 and 3.
We consider the $\Lambda$CDM model as a reference theory. In Tables 2 and 3, we compute the parameter $\Delta\left(AIC\right)$ for each constraint to assess the statistical preference of the models based on the given datasets. We note that the models $w_{\pm}^{I}\left(z\right),~{}w^{II}\left(z\right)$ and $w_{\pm}^{IV}\left(z\right)$ introduce only one additional degree of freedom compared to $\Lambda$CDM, whereas the models $w^{III}\left(z\right)$, $w^{V}\left(z\right)$, and $w^{VI}\left(z\right)$ have two extra free parameters.
@@ V.1 $SN+CC$ data
From the $SN+CC$ data, the $\Lambda$CDM model provides the maximum likelihood with the minimum value $\chi_{SN+CC}^{2}\left(\Lambda\right)=1419.6$. The models $w_{\pm}^{I}\left(z\right)$ fit the data with similar values for the physical parameters, with best-fit values $w_{a}^{I-}=-0.28$ and $w_{a}^{I+}=-0.05$, and a corresponding chi-square value of $\chi_{SN+CC}^{2}\left(w_{\pm}^{I}\right)=1419.8$. Thus, the difference in the Akaike Information Criterion (AIC) compared to $\Lambda$CDM is $\Delta\left(AIC\right)=2.2$.
For model $w^{II}\left(z\right)$, the best-fit value is $w_{a}=-0.40$, with a minimum chi-square value of $\chi_{SN+CC}^{2}\left(w^{II}\right)=1419.8$, leading to $\Delta\left(AIC\right)=2.2$. On the other hand, for model $w^{III}\left(z\right)$, we search for the best-fit values within the ranges $w_{a}\in\left[-30,10\right]$ and $w_{b}\in\left[-5,5\right]$. The analysis indicates that the two variables are degenerate, as previously discussed. Nevertheless, within this range, we find that $w_{a}<-12.5$ and $w_{b}=0.7$, with $\chi_{SN+CC}^{2}\left(w^{III}\right)=1419.6$ and $\Delta\left(AIC\right)=4.1$.
For the models of the second parametrization, we consider the parameter ranges $w_{a}\in(0,30]$ and $w_{b}\in[0,30]$ or $w_{b}\in[-10,10]$. However, this dataset is not sufficient to fully constrain the free parameters of this family of models. For models $w_{\pm}^{IV}\left(z\right)$, we find that $w_{a}=10.7$ and $w_{a}=10.4$, respectively, which are relatively large values, in agreement with the analysis presented previously in pant1 . The likelihood reaches its maximum value with $\chi_{SN+CC}^{2}\left(w_{\pm}^{IV}\right)=1419.5$, resulting in $\Delta\left(AIC\right)=1.9$.
@@
According to the AIC, for the SN+CC data, the models $\Lambda$CDM and $w_{\pm}^{IV}$ are statistically indistinguishable. Additionally, there is strong evidence that the models $w^{V}\left(z\right)$ and $w^{VI}\left(z\right)$ do not fit the data as well as $\Lambda$CDM.
@@
Figure 2: Evolution for the dark energy equation of state parameter $w_{\pm}^{I}\left(z\right),w^{II}\left(z\right)$ and $w^{III}\left(z\right)$ for the best fit values of Table 3.
Figure 3: Evolution for the dark energy equation of state parameter $w_{\pm}^{IV}\left(z\right),w^{V}\left(z\right)$ and $w^{VI}\left(z\right)$ for the best fit values of Table 3.
@@
$\Lambda$CDM $67.7_{-1.8}^{+1.8}$ $0.28_{-0.018}^{+0.018}$ $\nexists$ $\nexists$ $1419.6$ 00
$w_{+}^{I}\left(z\right)$ $67.5_{-1.7}^{+1.7}$ $0.29_{-0.022}^{+0.048}$ $-0.28_{-0.62}^{+0.62}$ $\nexists$ $1419.8$ $2.2$
$w_{-}^{I}\left(z\right)$ $67.6_{-1.7}^{+1.7}$ $0.28_{-0.031}^{+0.063}$ $-0.05_{-0.74}^{+0.74}$ $\nexists$ $1419.8$ $2.2$
$w^{II}\left(z\right)$ $67.4_{-1.8}^{+1.8}$ $0.30_{-0.039}^{+0.049}$ $-0.40_{-0.34}^{+1.1}$ $\nexists$ $1419.9$ $2.5$
$w^{III}\left(z\right)$ $67.7_{-1.7}^{+1.7}$ $0.28_{-0.027}^{+0.027}$ $<-12.5$ $0.7_{-2.4}^{+3.0}$ $1419.7$ $4.1$
$w_{+}^{IV}\left(z\right)$ $67.6_{-1.7}^{+1.7}$ $0.28_{-0.017}^{+0.020}$ $10.7_{-5.4}^{+.4}$ $\nexists$ $1419.5$ $1.9$
$w_{-}^{IV}\left(z\right)$ $67.7_{-1.7}^{+1.7}$ $0.29_{-0.019}^{+0.019}$ $10.4_{-5.5}^{+5.5}$ $\nexists$ $1419.5$ $1.9$
$w^{V}\left(z\right)$ $67.6_{-1.7}^{+1.7}$ $0.28_{-0.019}^{+0.022}$ $>5.7$ $>5.1$ $1419.5$ $3.9$
$w^{VI}\left(z\right)$ $67.8_{-1.7}^{+1.7}$ $0.28_{-0.018}^{+0.018}$ $>10.1$- $1419.5$ $4$
V.2 $SN+CC+BAO$ data
We perform a constraint on these models by additionally incorporating BAO data. The maximum likelihood value is again provided by the $\Lambda$CDM model, with $\chi_{\min}^{2}\left(\Lambda\right)=1435.4$. The models $w_{\pm}^{I}\left(z\right)$ fit the data with $\chi_{\min}^{2}\left(w_{+}^{I}\right)=1436.4$ and $\chi_{\min}^{2}\left(w_{-}^{I}\right)=1435.6$, respectively. However, due to the different number of free parameters, the corresponding $\Delta\left(AIC\right)$ values are $3.0$ and $2.2$, respectively.
On the other hand, model $w^{II}\left(z\right)$ fits the data with the same maximum likelihood value, that is, $\chi_{\min}^{2}\left(w^{II}\right)=1435.7$, leading to $\Delta\left(AIC\right)=2.1$. Moreover, for $w^{III}\left(z\right)$, we find $\chi_{\min}^{2}\left(w^{III}\right)=1433.6$, with $\Delta\left(AIC\right)=2.2$.

### v2

@@ Abstract
We investigate a dynamical reconstruction of the dark energy equation of state parameter by assuming that it satisfies a law of motion described by an autonomous second-order differential equation, with the limit of the cosmological constant as an equilibrium point. We determine the asymptotic solutions of this equation and use them to construct two families of parametric dark energy models, employing both linear and logarithmic parametrization with respect to the scale factor. We perform observational constraints by using the Supernova, the Cosmic Chronometers and the Baryon Acoustic Oscillations of DESI DR2. The constraint parameters are directly related with the initial value problem for the law of motion and its algebraic properties. The analysis shows that most of the models fit the observational data well with a preference to the models of the logarithmic parametrization. Furthermore, we introduce a new class of models as generalizations of the CPL model, for which the equilibrium point is a constant value rather than the cosmological constant. These models fit the data in a similar or better way to the CPL and the $\Lambda$CDM cosmological models.
@@ I Introduction
The determination of the equation of state parameter for the dark energy component can provide crucial insights into the nature of this exotic fluid source. Dynamical dark energy models with a time-varying equation of state parameter $w\left(a\right)$ have been introduced to explain cosmological observations. The Chevallier, Polarski and Linder cpl1 ; cpl1a (CPL) parametrization is the simplest varying dark energy model, introducing two degrees of freedom. The CPL model helps to determine how distinct a given dark energy model is from the cosmological constant cpl2 ; cpl3 ; cpl4 ; clp5 . Due to the limitations of the CPL model, various alternative dark energy models have been proposed over the past two decades. For more details, we refer the reader to par1 ; par2 ; par4 ; par5 ; par6 ; par7 ; exp1 ; exp2 ; osc1 ; osc2 ; osc3 ; osc4 ; osc5 ; pp1 and the references therein. The recent Baryon Acoustic Oscillations of DESI DR2  data indicate a preference for dynamical dark energy models over the $\Lambda$CDM model. The family of $w_{0}w_{a}$ des4 ; des5 ; des6 . This data set has been used to examined various theoretical models which lead to a dynamical dark energy bb1 ; bb2 ; bb3 ; bb4 ; bb5 ; bb6 ; bb7 ; bb8 ; bb9 ; bb10 ; bb11 ; bb12 ; bb13 ; bb14 .
@@ II Cosmological Framework
$\left(\dot{\rho}_{r}+4H\rho_{r}\right)+\dot{\rho}_{b}+\dot{\rho}_{dm}+\left(\dot{\rho}_{d}+3\left(1+w_{d}\right)H\rho_{d}\right)=0.$ (13)
@@
$\rho_{r}=\rho_{r,0}\left(\frac{a}{a_{0}}\right)^{-4},~{}\rho_{b}=\rho_{b,0}\left(\frac{a}{a_{0}}\right)^{-3},~{}\rho_{dm}=\rho_{dm,0}\left(\frac{a}{a_{0}}\right)^{-3}$ (14)
@@
$\Omega_{r0}+\Omega_{b0}+\Omega_{dm0}+\Omega_{b0}=1.$ (17)
@@
$H\left(a\right)=H_{0}\sqrt{\Omega_{r0}\left(\frac{a}{a_{0}}\right)^{-3}+\Omega_{b0}\left(\frac{a}{a_{0}}\right)^{-3}+\Omega_{dm0}\left(\frac{a}{a_{0}}\right)^{-3}+\Omega_{\Lambda 0}}.$ (18)
@@
$H\left(a\right)=H_{0}\sqrt{\Omega_{r0}\left(\frac{a}{a_{0}}\right)^{-3}+\Omega_{b0}\left(\frac{a}{a_{0}}\right)^{-3}+\Omega_{dm0}\left(\frac{a}{a_{0}}\right)^{-3}+\Omega_{d0}a^{-3\left(1+w_{0}\right)}}.$ (19)
@@ III Dynamical dark energy with $\Lambda$ as an Equilibrium Point
In this work we do not define explicitly the differential equation that the equation of state parameter satisfies. But we try to reconstruct the solution of the equation of state parameter near to the equilibrium point. For second-order theories, like quintessence, or phantom scalar field theories, the equation that $w_{d}\left(\chi\left(a\right)\right)$ satisfies is a first-order differential equation. On the other hand, for higher-order theories of gravity the order of differential equation is higher.
@@
where now $w_{a}=\operatorname{Re}\left(\Delta_{+}\right)=\operatorname{Re}\left(\Delta_{-}\right)$ and $w_{b}=\operatorname{Im}\left(\Delta_{+}\right)=\operatorname{Im}\left(\Delta_{-}\right)$. For a geometric interpretation of the stationary points we refer the reader in stro1 .
@@ III.3 Parametric space
$w_{+}^{I}\left(z\right)$ $w_{a}$ $-1+\left(e^{w_{a}Z}-1\right)$
$w_{-}^{I}\left(z\right)$ $w_{a}$ $-1-\left(e^{w_{a}Z}-1\right)$
@@ IV Observational Data
  • (SN) The Pantheon+ dataset [^1] https://github.com/PantheonPlusSH0ES/DataRelease comprises 1701 light curves of 1550 spectroscopically confirmed supernova events within the range $10^{-3}<z<2.27~{}$pant . The data provides the distance modulus $\mu^{obs}~{}$at observed redshifts $z$. The theoretical distance modulus is constructed as follows $\mu^{th}=5\log D_{L}+25,$ where $D_{L}$ is the luminosity distance defined by the Hubble functions from the expression $D_{L}=c\left(1+z\right)\int\frac{dz}{H\left(z\right)}$.
  • (CC) For the Cosmic Chronometers we use 31 direct measurements of the Hubble parameter across redshifts in the range [^2] https://github.com/Ahmadmehrabi/Cosmic_chronometer_data $0.09\leq z\leq 1.965~{}$ as summarized in Table 1 in sunny .
  • (BAO) The BAO data are provided by the SDSS Galaxy Consensus, quasars and Lyman-$\alpha$ forests bbn0 and the data from the DESI DR1 collaboration des1 ; des2 ; des3 where we refer as BAO1 and the DESI DR2 collaboration des4 ; des5 ; des6 which we refer as BAO2. These data sets provide observation values of the $\frac{D_{M}}{r_{s}},~{}\frac{D_{V}}{r_{s}}$, $\frac{D_{H}}{r_{s}}$ where $r_{s}$ is the sound horizon at the drag epoch, $D_{M}$ is the comoving angular distance $D_{M}=\left(1+z\right)^{-1}D_{L}$ ; $D_{V}$ is the volume averaged distance $D_{V}^{3}=cD_{L}\frac{z}{H\left(z\right)}$ and $D_{H}$ is the Hubble distance $D_{H}=\frac{c}{H}$.
@@ IV.1 Methodology
In order to obtain the optimal values for the cosmological parameters we employ the Bayesian inference COBAYA [^3] https://cobaya.readthedocs.io/ cob1 ; cob2 with a custom theory and the MCMC sampler. From the results of the Planck 2018 collaborations pl1n we consider the energy density for the radiation to be $\Omega_{r0}=4.15~{}10^{-5}$. Thus, we constrain our proposed models for the parameters $\left\{H_{0},\Omega_{m0},\mathbf{w,}r_{drag}\right\},~{}$where $\mathbf{w}$ describes the vector of parameters within the equation of state parameters, as they are given in Table 1 and $\Omega_{m0}=\Omega_{dm0}+\Omega_{b0}$.
For all the runs we selected the priors $H_{0}\in\left[60,75\right]$,${}\Omega_{m0}\in\left[0.2,0.45\right]$ and $r_{drag}\in\left[130,160\right]$.
@@ V Results
We constrain the dynamical dark energy models using two sets of data: (I) Supernova and Cosmic Chronometers ($SN+CC$) and (II) Supernova, Cosmic Chronometers and Baryonic Acoustic Oscillations of DESI DR1 ($SN+CC+BAO_{1}$) and (III) Supernova, Cosmic Chronometers and Baryonic Acoustic Oscillations of DESI DR2 ($SN+CC+BAO_{2}$). We determine the free parameters along with their corresponding $1\sigma$ errors by locating the maximum of the likelihood functions. The results for the two tests are presented in Tables 2, 3 and 4.
We consider the $\Lambda$CDM model as a reference theory. In Tables 2, 3 and 4, we compute the parameter $\Delta\left(AIC\right)$ for each constraint to assess the statistical preference of the models based on the given datasets. We note that the models $w_{\pm}^{I}\left(z\right),~{}w^{II}\left(z\right)$ and $w_{\pm}^{IV}\left(z\right)$ introduce only one additional degree of freedom compared to $\Lambda$CDM, whereas the models $w^{III}\left(z\right)$, $w^{V}\left(z\right)$, and $w^{VI}\left(z\right)$ have two extra free parameters. In Figs. 2 and 3 we present the qualitative evolution of the dynamical dark energy parameters.
@@ V.1 $SN+CC$ data
From the $SN+CC$ data, the $\Lambda$CDM model provides the maximum likelihood with the minimum value $\chi_{SN+CC}^{2}\left(\Lambda\right)=1419.5$. The models $w_{\pm}^{I}\left(z\right)$ fit the data with similar values for the physical parameters, with best-fit values $w_{a}^{I-}=-0.24$ and $w_{a}^{I+}=-0.05$, and a corresponding chi-square value of $\chi_{SN+CC}^{2}\left(w_{+}^{I}\right)=1419.8$ and $\chi_{SN+CC}^{2}\left(w_{+}^{I}\right)=1419.9$. Thus, the difference in the Akaike Information Criterion (AIC) compared to $\Lambda$CDM are $\Delta\left(AIC\right)=2.3$ and $2.4$ respectively.
For model $w^{II}\left(z\right)$, the best-fit value is $w_{a}=-0.37$, with a minimum chi-square value of $\chi_{SN+CC}^{2}\left(w^{II}\right)=1419.9$, leading to $\Delta\left(AIC\right)=2.4$. On the other hand, for model $w^{III}\left(z\right)$, we search for the best-fit values within the ranges $w_{a}\in\left[-30,10\right]$ and $w_{b}\in\left[-5,5\right]$. The analysis indicates that the two variables are degenerate, as previously discussed. Nevertheless, within this range, we find that $w_{a}<-12.6$ and $w_{b}=1.1$, with $\chi_{SN+CC}^{2}\left(w^{III}\right)=1419.7$ and $\Delta\left(AIC\right)=4.2$.
[…]
@@
According to the AIC, for the SN+CC data, the models $\Lambda$CDM and $w_{\pm}^{IV}$ are statistically indistinguishable. Additionally, there is strong evidence that the models $w^{III}\left(z\right)$, $w^{V}\left(z\right)$ and $w^{VI}\left(z\right)$ do not fit the data as well as $\Lambda$CDM.
@@
Figure 2: Evolution for the dark energy equation of state parameter $w_{\pm}^{I}\left(z\right),w^{II}\left(z\right)$ and $w^{III}\left(z\right)$.
Figure 3: Evolution for the dark energy equation of state parameter $w_{\pm}^{IV}\left(z\right),w^{V}\left(z\right)$ and $w^{VI}\left(z\right)$.
@@
$\Lambda$CDM $67.5_{-1.9}^{+1.6}$ $0.331^{+0.017}$ $\nexists$ $\nexists$ $1419.5$ 00
$w_{+}^{I}\left(z\right)$ $67.5_{-1.7}^{+1.7}$ $0.337_{-0.025}^{+0.049}$ $-0.24_{-0.66}^{+0.76}$ $\nexists$ $1419.8$ $2.3$
$w_{-}^{I}\left(z\right)$ $67.7_{-1.8}^{+1.8}$ $0.324_{-0.032}^{+0.063}$ $-0.05_{-0.74}^{+0.74}$ $\nexists$ $1419.9$ $2.4$
$w^{II}\left(z\right)$ $67.5_{-1.8}^{+1.8}$ $0.345_{-0.038}^{+0.052}$ $-0.37_{-0.65}^{+1.1}$ $\nexists$ $1419.9$ $2.4$
$w^{III}\left(z\right)$ $67.7_{-1.9}^{+1.7}$ $0.327_{-0.029}^{+0.029}$ $<-12.6$ $1.1_{-2.6}^{+2.6}$ $1419.7$ $4.2$
$w_{+}^{IV}\left(z\right)$ $67.6_{-1.8}^{+1.8}$ $0.328_{-0.017}^{+0.022}$ $>6.6$ $\nexists$ $1419.7$ $2.2$
$w_{-}^{IV}\left(z\right)$ $67.7_{-1.7}^{+1.7}$ $0.333_{-0.020}^{+0.018}$ $10.1_{-8.3}^{+6.4}$ $\nexists$ $1419.5$ $2.0$
$w^{V}\left(z\right)$ $67.7_{-1.7}^{+1.7}$ $0.331_{-0.024}^{+0.024}$ $>5.63$ $>4.92$ $1419.5$ $4.0$
$w^{VI}\left(z\right)$ $67.8_{-1.7}^{+1.7}$ $0.326_{-0.010}^{+0.019}$ $15.7_{-8.3}^{+8.3}$- $1419.2$ $3.7$
V.2 DESI  DR1: $SN+CC+BAO_{1}$ data
We perform a constraint on these models by additionally incorporating BAO1 data. The maximum likelihood value is again provided by the $\Lambda$CDM model, with $\chi_{\min}^{2}\left(\Lambda\right)=1435.2$. The models $w_{\pm}^{I}\left(z\right)$ fit the data with $\chi_{\min}^{2}\left(w_{+}^{I}\right)=1435.8$ and $\chi_{\min}^{2}\left(w_{-}^{I}\right)=1435.4$, respectively. However, due to the different number of free parameters, the corresponding $\Delta\left(AIC\right)$ values are $2.6$ and $2.2$, respectively.
On the other hand, model $w^{II}\left(z\right)$ fits the data with the same maximum likelihood value, that is, $\chi_{\min}^{2}\left(w^{II}\right)=1435.4$, leading to $\Delta\left(AIC\right)=2.2$. Moreover, for $w^{III}\left(z\right)$, we find $\chi_{\min}^{2}\left(w^{III}\right)=1433.5$, with $\Delta\left(AIC\right)=2.3$.